what is 9+(3)-5+(-4)
Found 2 solutions by Mr.Math, sofiyac:
Answer by Mr.Math(29)

9+(3)-5+(-4)= 9 + 3 - 5 - 4 = 3

Answer by sofiyac(983)

9+(3)-5+(-4)=9+3-5-4=12-5-4=7-4=3